# A Systematic Literature Review on Detecting Software Vulnerabilities with Large Language Models

### v1

@@ 3. Methodology
We followed established guidelines by Kitchenham et al. (Kitchenham and Charters, 2007; Petersen et al., 2015) for conducting SLRs in software engineering research. We investigate the following research questions RQ1–RQ4, cf. Table 2: RQ1 How is the vulnerability detection task formulated and approached in LLM-based systems? RQ2 How is input, particularly the semantics of vulnerabilities, represented and encoded for LLMs? RQ3 What are the predominant techniques for adapting LLMs to the task of software vulnerability detection? RQ4 What datasets are used to evaluate LLM-based vulnerability detection approaches, and how do dataset choices impact comparability across studies? A joint investigation of RQ1–RQ4 is crucial to obtain a holistic understanding of LLM-based vulnerability detection research, from input representations to system design and evaluation, and to identify key levers to enhance robustness, generalization, and comparability. We conducted the SLR in three phases, i.e., (1) literature search, (2) study selection, and (3) study analysis, as detailed in the following.
| Research Question | Objective | Discussion | |
| RQ1 | How is the vulnerability detection task formulated and approached in LLM-based systems? | This question focuses on how the detection task is formulated, and how this formulation shapes system architecture. | Section 4.1 and Section 4.3 |
| RQ2 | How is input, particularly the semantics of vulnerabilities, represented and encoded for LLMs? | This question investigates how code and vulnerability-related semantics are formatted and passed to the model, with a focus on how effectively vulnerabilities are represented. | Section 4.2 |
| RQ3 | What are the predominant techniques for adapting LLMs to the task of software vulnerability detection? | Here, the focus lies on identifying and categorizing state-of-the-art adaptation strategies, including prompt engineering styles, fine-tuning techniques, and advanced training paradigms. | Section 4.4 |
| RQ4 | What datasets are used to evaluate LLM-based vulnerability detection approaches, and how do dataset choices impact comparability across studies? | This question explores the characteristics of datasets used in current research (such as type, size, and label quality), how they affect model generalization, benchmarking, and the overall comparability of experimental results across studies. | Section 5 |
@@ 3.1. Literature Search
To identify relevant studies, we conducted a systematic literature search across three major databases: IEEE Xplore, ACM Digital Library, and arXiv.
We selected IEEE Xplore and the ACM Digital Library as they host peer-reviewed publications from conferences and journals.
We further include arXiv for preprints to capture ongoing research that is still in the submission process, given the rapidly evolving nature of this field.
We manually identified a set of relevant primary studies in our prior publication (Kaniewski et al., 2024), which we used to derive keywords for an initial search string.
We refined this search string using keywords from related surveys (Hou et al., 2024; Zhang et al., 2024a).
We searched the selected databases for studies whose titles or abstracts matched this search string.
Before the cutoff date in July, the database search yielded a total of 552 studies from IEEE Xplore, 344 studies from ACM, and 2287 studies from arXiv.
For reproducibility, the final search string is as follows:
Keywords: (LLM OR LLMs OR Large Language Model OR Large Language Models OR pre-train* OR GPT* OR ChatGPT OR T5 OR LLaMA* OR Codex OR BERT) AND ((vulnerability OR vulnerabilities OR CVE OR CWE) AND (detect* OR identif* OR classif* OR analy* OR discover* OR assess*) AND (software OR program OR code))
@@ 3.2. Study Selection
Following the initial retrieval, we applied a selection process to filter irrelevant studies. We reviewed the title and abstract of each study and screened it to assess its relevance based on predefined inclusion (✓) and exclusion (✗) criteria:
Studies that use and evaluate LLMs for software vulnerability detection, introducing novel methods or systems.
Studies that do not use LLMs to actively detect vulnerabilities but solely to work on outputs produced by other (static analysis) tools.
Studies that mention LLMs only in passing or use them merely as baselines.
Studies that focus on domains outside source code vulnerability detection, such as binary code, smart contracts, network analysis, or malware detection.
Studies not written in English.
Studies not accessible in full text.
Studies identified as duplicates or minor variations of multi- or extended-version papers from the same authors.
Studies classified as reviews, surveys, theses, abstracts, posters, demonstrations, or tutorials.
Given the inclusion of preprints from arXiv, we conducted a manual quality check to determine study eligibility. We assessed whether the studies provided a clear contribution, adequately described the proposed workflow and implementation, outlined the evaluation setup (including datasets, baselines, and metrics), and presented coherent findings that support their central claims.
Finally, we performed backward and forward snowballing. We manually reviewed the reference lists of the selected studies (backward snowballing) and used citation tracking via Google Scholar to identify and evaluate additional papers citing the included studies (forward snowballing).
@@ 3.3. Study Analysis
Following the selection process, we identified 227 papers that met our selection criteria. The earliest included study dates back to 2020. The vast majority of publications (106 studies, 46.7%) were published in 2024, followed by 54 studies in 2025, underscoring the rapidly growing interest in and significance of LLM-based vulnerability detection. Notably, 83 (36.6%) of the selected studies were sourced from arXiv, reflecting both the fast-paced development of this field and the prevalence of recent works still undergoing peer review.
@@
During the full-text review, we extracted key data points, recording how the vulnerability detection task was defined, which LLMs were used, and what adaptation techniques were employed. Additionally, we extracted information on input representation formats, datasets used for fine-tuning and evaluation, targeted vulnerability types (e.g., specific CWEs), and evaluation metrics. This structured analysis enabled us to derive a comprehensive taxonomy of LLM-based software vulnerability detection approaches, address the research questions, and develop a deeper understanding of current trends, open challenges, and emerging directions in LLM-based software vulnerability detection.
4. Vulnerability Detection with Language Models
@@
Recent research has adopted a variety of techniques for detecting software vulnerabilities using LLMs. In this chapter, we present a comprehensive taxonomy that categorizes the different ways LLMs are applied to vulnerability detection. Specifically, we examine task formulations, input representations, system architectures, and adaptation techniques. The overall taxonomy is illustrated in Figure 1. We discuss each category on the basis of the surveyed studies in the following subsections.
Taxonomy showing categories and attribute values for LLM-based vulnerability detection studies.
4.1. Detection Task
LLM-based vulnerability detection can be framed in multiple ways, depending on how the task is formulated and what objectives are prioritized beyond detection. This section outlines and discusses the most common formulations, particularly as classification tasks, and additional objectives.
4.1.1. Task Formulation
@@
Vulnerability detection is most commonly formulated as a classification problem, differentiating binary classification, vulnerability-specific classification, and multi-class classification (Shereen et al., 2024). Binary Classification determines whether a given code contains a security vulnerability or not (’Yes’/’No’). Vulnerability-Specific Classification, as a refinement of the binary formulation, determines whether a given code contains a specific vulnerability, represented, e.g., by CWE-ID. Multi-Class Classification determines which specific type of vulnerability the given code contains, often using CWE-IDs as class labels. This formulation may use a given pre-defined list of vulnerability types in a prompt engineering setting.
The choice of task formulation depends on factors such as the application domain, the availability and quality of labeled data, and the surrounding tooling and infrastructure. The surveyed studies use one or multiple of these formulations. With 156 studies, binary classification remains the most prevalent task formulation, cf. (Shereen et al., 2024), followed by multi-class classification with 90 studies. Only 27 studies (Atiiq et al., 2024b; Bae et al., 2024; Cao et al., 2024; Ceka et al., 2025; Dolcetti et al., 2025; Gong et al., 2025; Guo et al., 2025; Khare et al., 2025; Li et al., 2025d; Mao et al., 2025; Mohajer et al., 2024; Nong et al., 2024; Szabó and Bilicki, 2023; Ullah et al., 2024; Yang et al., 2025a; Zibaeirad and Vieira, 2024; Li et al., 2025a; Mächtle et al., 2025; Mou et al., 2025; Munson et al., 2025; Seo et al., 2025; Tian et al., 2025; Wang et al., 2025a; Zhang et al., 2024c; Zhou et al., 2025b; Wang et al., 2025b; Li et al., 2025c) adopt a vulnerability-specific classification. Notably, Zibaeirad and Vieira (Zibaeirad and Vieira, 2024) systematically apply and compare all three task formulations using different prompts.
@@
Binary classification simplifies the detection task as it is reduced to a decision of whether a code sample is vulnerable or not. While this approach facilitates model training and deployment, it lacks fine-grained insight into the specific type or cause of the vulnerability, limiting its use in real-world vulnerability remediation.
@@
Vulnerability-specific classification is more targeted, which is valuable in contexts involving high-impact CWEs or regulatory compliance. However, it depends on curated datasets and may struggle to generalize to related vulnerability types not present in the training data.
@@
Multi-class classification provides the most detailed perspective by distinguishing among various types of vulnerabilities. This formulation aligns more closely with real-world vulnerability remediation, where identifying the exact nature of a vulnerability is essential for vulnerability patching. However, it introduces additional complexity in both modeling and evaluation. Its effectiveness also depends heavily on high-quality, CWE-labeled, and well-distributed datasets.
@@
4.1.2. Additional Objectives
While classification is the primary formulation for vulnerability detection, it offers limited insight into the exact location or root cause of vulnerabilities. Consequently, several studies have extended the scope to incorporate additional objectives that align more closely with real-world remediation workflows. With the growing adoption of LLMs for Code Generation, several studies (Ambati et al., 2024; Dai et al., 2025; Dolcetti et al., 2025; Gong et al., 2025; Liu et al., 2024c; Nunez et al., 2024; Wang et al., 2023a; Yu et al., 2024; Mou et al., 2025; Li et al., 2025c) investigated the vulnerabilities present in LLM-generated code, often using the same LLM for both generation and detection (self-evaluation capabilities). Other studies extend the vulnerability detection task to include additional stages in the vulnerability assessment pipeline, such as Vulnerability Localization (Bahaa et al., 2024; Bakhshandeh et al., 2023; Curto et al., 2024b; Ding et al., 2023; Du et al., 2024; Fu et al., 2024; Fu and Tantithamthavorn, 2022; Islam et al., 2024a; Le et al., 2024; Li et al., 2024; Liu et al., 2024a; Mao et al., 2025; Mohajer et al., 2024; Pelofske et al., 2024; Shashwat et al., 2024; Sovrano et al., 2025; Steenhoek et al., 2025b, a; Tamberg and Bahsi, 2025; Widyasari et al., 2024; Yin et al., 2024; Yu et al., 2025; Ahmed et al., 2025; Halder et al., 2025; Mächtle et al., 2025; Mou et al., 2025; Peng et al., 2023; Wen et al., 2025; Chen et al., 2025), Severity Estimation (Fu et al., 2024, 2023b; Torkamani et al., 2025; Liu et al., 2025b; Mahmud et al., 2025; Widyasari et al., 2025), Vulnerability Repair (Ambati et al., 2024; Dolcetti et al., 2025; Du et al., 2025; Espinha Gasiba et al., 2024; Ezenwoye et al., 2024; Fu et al., 2024, 2023b; Gong et al., 2025; Islam et al., 2024b; Khan et al., 2025; Kouliaridis et al., 2024; Liu et al., 2024c; Mao et al., 2025; Noever, 2023; Nunez et al., 2024; Steenhoek et al., 2025b; Tamberg and Bahsi, 2025; Wu et al., 2023; Yang et al., 2024a; Yu et al., 2025; Zibaeirad and Vieira, 2024; Liu et al., 2025b; Mou et al., 2025; Seo et al., 2025; Wang et al., 2025b), or Security Testing (Nunez et al., 2024; Liu et al., 2025b; Wang et al., 2025a), e.g., validating the results by running the security tests generated (Wang et al., 2025a). In addition to classification outputs, many approaches prompt for natural language descriptions of the identified vulnerabilities (extending the classification into a generation task), typically referring to the CWE-type in multi-class classification. With the emergence of more advanced reasoning models and structured prompting techniques, studies (Cipollone et al., 2025; Du et al., 2024, 2025; Ingemann Tuffveson Jensen et al., 2024; Islam et al., 2024a; Li et al., 2025d; Liu et al., 2024c; Mao et al., 2025; Mohajer et al., 2024; Ni et al., 2023b; Shashwat et al., 2024; Steenhoek et al., 2025a; Sun et al., 2025; Tamberg and Bahsi, 2025; Ullah et al., 2024; Wu et al., 2023; Yang et al., 2025c; Yildiz et al., 2025; Yu et al., 2025; Zibaeirad and Vieira, 2025; Ahmed et al., 2025; Li et al., 2025a, e; Lingxiang et al., 2025; Liu et al., 2025b; Mächtle et al., 2025; Mou et al., 2025; Seo et al., 2025; Wang et al., 2025a; Wen et al., 2025; Widyasari et al., 2025; Zhou et al., 2025b; Wang et al., 2025b; Weyssow et al., 2025) shift towards Reasoning about root causes of vulnerabilities. For example, Steenhoek et al. (Steenhoek et al., 2025a) perform an error analysis on LLM responses, demonstrating that the models struggle to reason about the code semantics relevant to identifying vulnerabilities, especially subtle semantic differences caused by small textual changes. Wen et al. (Wen et al., 2025) propose forward and backward reasoning, where the forward reasoning aims to deduce the causes of vulnerabilities, while the backward reasoning seeks to understand the code changes in the fixes. Their results show that LLMs can reason about vulnerabilities but often lack the domain-specific knowledge required to effectively invoke such reasoning.

### v2

@@ 3. Methodology
We followed established guidelines by Kitchenham et al. (Kitchenham and Charters, 2007; Petersen et al., 2015) for conducting SLRs in software engineering, i.e., defining research questions, conducting the search, defining selection criteria, and analyzing the selected studies. In the following, we outline the research focus of this SLR as well as the three phases of literature search, study selection, and study analysis, as visualized in Figure 1.
3.1. Research Focus
In this SLR, we address the following research questions RQ1–RQ5:
RQ1
How is the vulnerability detection task formulated and approached in LLM-based systems? Traditional deep-learning-based approaches predominantly frame the detection task as binary classification (vulnerable vs. non-vulnerable) (Shereen et al., 2024). While this formulation simplifies evaluation, it falls short in reflecting real-world software engineering workflows, where developers additionally require information on, e.g., vulnerability types and root causes. As LLMs are capable of producing structured reasoning steps and conceptual explanations, understanding whether LLM-based approaches move beyond this limited formulation is essential. We investigate how existing LLM-based approaches formulate the task, i.e., do they maintain the binary framing or do they extend or reformulate it, e.g., as multi-label vulnerability type prediction or vulnerability report generation (cf. Section 4.1).
RQ2
How is input, particularly the semantics of vulnerabilities, represented and encoded for LLMs? Existing machine learning and deep learning approaches convert code into suitable representations to capture semantics, for example, by constructing graph structures for GNNs or vectorizing code embeddings for CNNs (Shiri Harzevili et al., 2024; Shereen et al., 2024). For LLMs with their world knowledge, it remains unclear whether code alone is sufficient to capture specific vulnerability semantics. Investigating how studies represent input and whether they incorporate auxiliary information is crucial to understanding what information LLMs can leverage for effective detection (cf. Section 4.2).
RQ3
What are the predominant techniques for adapting LLMs to the task of software vulnerability detection? LLMs can be adapted via prompting, fine-tuning, or pre-training. Each technique has different implications for data requirements, computational costs, and practical deployment. Since these constraints affect the feasibility of integrating LLM-based methods into software engineering pipelines, this question investigates which adaptation techniques are used and how well they align with practical constraints (cf. Section 4.4.2).
RQ4
What datasets are used to train, tune, or evaluate LLM-based vulnerability detection approaches, and how does dataset selection impact applicability? Dataset selection influences the types and variety of vulnerabilities under detection. In terms of applicability, synthetic or oversimplified datasets lead to models with poor generalization. Therefore, understanding the characteristics (such as covered vulnerability types, size, and label quality) of the datasets used in current research is essential to enable reliable benchmarking and assess model generalization to real-world applications (cf. Section 5).
RQ5
How does dataset selection impact comparability across studies? Dataset selection influences whether results are comparable across studies. Inconsistent evaluation setups as well as not adhering to open science practices make it impossible to compare study results. Identifying commonly used datasets as part of evaluation setups enables researchers to select datasets to improve comparability across studies (cf. Section 5.3).
A joint investigation of RQ1-RQ5 is crucial to obtain a holistic understanding of LLM-based vulnerability detection research, from input representations to system design and evaluation, and to identify key levers to enhance robustness, practicability, and comparability.
@@ 3.2. Literature Search
To identify relevant studies, we conducted a systematic literature search across three major databases: IEEE Xplore, ACM Digital Library, and arXiv. We selected IEEE Xplore and the ACM Digital Library, as they host peer-reviewed publications from conferences and journals. We further include arXiv for preprints to capture ongoing research that is still in the submission process, given the rapidly evolving nature of this field.
The search string construction followed a two-step process.
First, we manually identified a set of relevant primary studies in our prior publication (Kaniewski et al., 2024).
We used these studies as a starting point to derive vulnerability detection-related terms. Second, to ensure systematic coverage of LLM-related terminology, we incorporated keywords used in related SLRs (Hou et al., 2024; Zhang et al., 2024a).
Specifically, we considered the keywords LLM, Large Language Model, Pre-trained, Pre-training, BERT, Codex, GPT, T5, ChatGPT, and GPT-*. We iteratively refined the search string by adding observed major model families and used wildcard extensions to ensure systematic coverage. The final search string is as follows:
Search String: (LLM OR LLMs OR Large Language Model OR Large Language Models OR pre-train* OR GPT* OR ChatGPT OR T5* OR CodeT5* OR Llama* OR CodeLlama* OR Codex OR BERT OR CodeBERT OR GraphCodeBERT OR Bard OR Claude OR Gemma* OR DeepSeek* OR Gemini* OR Mistral* OR Mixtral* OR Phi* or Qwen* OR CodeQwen* OR WizardCoder*) AND ((vulnerability OR vulnerabilities OR CVE OR CWE) AND (detect* OR identif* OR classif* OR analy* OR discover* OR assess*) AND (software OR program OR code))
@@ 3.3. Study Selection
Figure 1 illustrates the study selection pipeline used in this SLR. Using the final search string, we ran an automated search across the selected databases to identify studies whose titles or abstracts matched the specified keywords. We carried out the initial search in December 2024. After the initial crawl, we continuously integrated newly published material through database-specific search alerts and scholar.inbox (Flicke et al., 2025), ensuring that further relevant studies were considered and included.
Following this automated search, we applied a selection process to filter out irrelevant studies based on predefined inclusion (✓) and exclusion (✗) criteria, cf. Table 2. We retrieved each accessible study written in English and first filtered by title and abstract before performing a full text screen to assess if it is in scope. Specifically, we focus on studies that apply LLMs to actively detect software vulnerabilities, rather than applying LLMs to work with outputs of other detection tools, e.g., SAST tools. Further, we specifically focus on software vulnerability detection rather than detection in binary code, smart contracts, network logs, or detection of malware. We further exclude any duplicates or minor variations of multi- or extended-version papers as well as studies classified as reviews, surveys, theses, abstracts, posters, demonstrations, or tutorials.
Given the inclusion of preprints from arXiv, we conducted a manual quality assessment to determine study eligibility. We assessed whether the studies provided a clear contribution, adequately described the proposed workflow and implementation, outlined the evaluation setup (including datasets, baselines, and metrics), and presented coherent findings that support their central claims.
After filtering, we performed backward and forward snowballing. We manually reviewed the reference lists of the selected studies (backward snowballing) and used citation tracking via Google Scholar to identify additional papers citing the included studies (forward snowballing). Studies retrieved through snowballing also pass through the pipeline of inclusion and exclusion criteria.
Finally, at the time of writing this SLR, we cover 263 studies. Further studies added through the continuous search will be made available as part of the maintained repository, cf. (Kaniewski et al., 2025).
todo.
| Inclusion Criteria |
| ✓ Studies with title or abstract matching the specified search string. |
| ✓ Studies that use and evaluate LLMs for software vulnerability detection, introducing novel techniques or systems. |
| Exclusion Criteria |
| ✗ Studies not written in English. |
| ✗ Studies not accessible in full text. |
| ✗ Studies classified as reviews, surveys, theses, abstracts, posters, demonstrations, or tutorials. |
| ✗ Studies that focus on domains outside source code vulnerability detection, such as binary code, smart contracts, network analysis, or malware detection. |
| ✗ Studies that do not use LLMs to actively detect vulnerabilities but solely to work on outputs produced by other tools. |
| ✗ Studies that mention LLMs only in passing or use them merely as baselines. |
| ✗ Studies identified as duplicates or minor variations of multi- or extended-version papers from the same authors. |
@@ 3.4. Study Analysis
Following the selection process, we identified 263 papers that met the selection criteria. During the full-text review of each selected study, we extracted key data points, recording how the vulnerability detection task was defined, which LLMs were used, and what adaptation techniques were employed. Additionally, we extracted information on input representation formats, datasets used for fine-tuning and evaluation, targeted vulnerability types (e.g., specific CWEs), and evaluation metrics. This structured analysis enabled us to derive a comprehensive taxonomy of LLM-based software vulnerability detection approaches, address the research questions, and develop a deeper understanding of current trends, open challenges, and emerging directions in LLM-based software vulnerability detection.
The earliest included study dates back to 2020 (cf. Fig 2 (a)). Since then, we can observe a rapid increase in LLM-based vulnerability detection studies. Notably, 78 (29.7%) of the selected studies were sourced from arXiv, cf. Fig 2 (b), reflecting both the fast-paced development of this field and the prevalence of recent works still undergoing peer review. For the peer-reviewed studies, Fig 2 (c) visualizes the distribution of the selected studies across venue tiers. In addition, Table 3 lists the respective A*/A conferences as well as Q1 journals with their study counts.
| Type | Venue (Count) |
| Conf. A* | ICSE (13), EMNLP☼ (4), USENIX Security† (3), ASE (3), ESEC/FSE (2), INFOCOM◦ (2), ACL◦ (2), NDSS† (1), SP† (1), ICLR☼ (1) |
| Conf. A | ISSRE (3), ESORICS† (3), ACSAC† (2), ISSTA (2), SANER (2), MSR (2), DSN◦ (1), EASE (1), EuroS&P† (1), ASIACCS† (1), ICPC (1), ICSME (1), ICST (1), NAACL◦ (1), RAID† (1) |
| Jour. Q1 | TSE (7), IEEE Access◦ (4), COSE† (4), EMSE (4), TOSEM (3), JISA† (3), JSS (3), IST (1), Info. Fusion◦ (1), Cluster Computing◦ (1), TDSC† (1), Neural Comput. Appl.◦ (1) |
4. Vulnerability Detection with LLMs
@@
Recent research has adopted a variety of techniques for detecting software vulnerabilities using LLMs. In this chapter, we present a comprehensive taxonomy that categorizes the different ways LLMs are applied to vulnerability detection. Specifically, we examine task formulations, input representations, system architectures, and techniques. The overall taxonomy is illustrated in Figure 3. We discuss each category on the basis of the surveyed studies in the following subsections.
4.1. Task Formulation (F)
Software vulnerability detection can be framed in multiple ways, depending on how the task is formulated and what objectives are prioritized beyond detection (e.g., explanation). With respect to the objectives, the integration of LLMs into vulnerability detection introduces unique advantages over traditional approaches. Unlike rule-based approaches, which require definitions of vulnerable patterns, the language understanding capabilities of LLMs allow for highly flexible goal setting; detection objectives can be refined or expanded, e.g., adding reasoning or vulnerability localization, simply by adjusting natural language prompts. LLMs further have extensive pre-trained world knowledge, such as software architecture patterns, programming language-specific constructs, and broad cybersecurity concepts. Such domain knowledge is difficult to encode explicitly in traditional tools or classical ML approaches, but serves as a versatile starting point for LLM-based detection. In the following, we outline and discuss the most common detection task formulations used with LLMs in the surveyed studies: classification and generation tasks.
4.1.1. Classification (F1)
@@
Vulnerability detection is most commonly formulated as a classification problem, differentiating binary and multi-class forms. Binary Classification (F1.1) determines whether a given code contains a security vulnerability or not (’Yes’/’No’). Vulnerability-Specific Classification (F1.1.1) (Shereen et al., 2024), as a refinement of the binary formulation, determines whether a given code contains a specific vulnerability, represented, e.g., by CWE-ID. Multi-Class Classification (F1.2) determines which specific type of vulnerability the given code contains (type prediction), often using CWE-IDs as class labels. This formulation may use a given pre-defined list of vulnerability types in a prompt engineering setting. Similarly, Multi-Label Classification (F1.3) aims to determine any vulnerability types present in the code, recognizing that a code snippet may contain multiple distinct vulnerabilities simultaneously.
@@
Binary classification (F1.1, 195 studies) simplifies the detection task as it is reduced to a decision of whether a code sample is vulnerable or not. While this approach facilitates model training and deployment, it lacks fine-grained insight into the specific type or cause of the vulnerability, limiting its use in real-world vulnerability remediation.
@@
Vulnerability-specific classification (F1.1.1, 20 studies) is more targeted, which is valuable in contexts with high-impact CWEs or regulatory compliance. However, it depends on curated datasets and may struggle to generalize to related vulnerability types not present in the training data.
@@
Multi-class (F1.2, 61 studies) and multi-label (F1.3, 23 studies) classification provide the most detailed perspective by distinguishing among various types of vulnerabilities. These formulations align more closely with real-world vulnerability remediation, where identifying the exact nature of a vulnerability is essential for vulnerability patching. However, these formulations introduce additional complexity in both modeling and evaluation. Their effectiveness also depends heavily on high-quality, CWE-labeled, and well-distributed datasets.
@@
4.1.2. Generation (F2)
While classification is the primary formulation, it offers limited explanations and insights into the root causes of vulnerabilities. Therefore, several studies extend the scope of classification and leverage the language understanding and generation capabilities of LLMs to generate additional outputs. This generative approach takes advantage of the model’s flexibility, allowing researchers to request diverse detection outputs, such as reasoning chains or fix suggestions, through simple natural language instructions. We categorize generative tasks as (i) vulnerability description, (ii) root cause reasoning, and (iii) structured vulnerability report.
Vulnerability Description (F2.1) focuses on generating a high-level natural language summary or description of the specific detected vulnerability, providing developers with its context (Chan et al., 2023; Cipollone et al., 2025; Ezenwoye et al., 2024; Fu et al., 2023a; Gajjar et al., 2025a; Ingemann Tuffveson Jensen et al., 2024; Islam et al., 2023; Kouliaridis et al., 2024; Lin and Mohaisen, 2024; Li et al., 2025d, c; Mohajer et al., 2024; Safdar et al., 2025a; Shu et al., 2025; Sovrano et al., 2025; Steenhoek et al., 2025b; Tamberg and Bahsi, 2025; Wang et al., 2025g, d; Yıldırım et al., 2024; Yildiz et al., 2025).
Root Cause Reasoning (F2.2) moves beyond high-level descriptions to provide a code-level justification or reasoning of the specific vulnerability and its root cause (Ahmed et al., 2025; Dolcetti et al., 2025; Du et al., 2025a; Espinha Gasiba et al., 2024; Hairan and Şahman, 2025; Islam et al., 2024b; Jararweh et al., 2025; Keltek et al., 2024; Khare et al., 2025; Kholoosi et al., 2024; Li et al., 2025g, i, f, h; Liu et al., 2023, 2024b; Li et al., 2025c; Mao et al., 2025; Mathews et al., 2024; Mhatre et al., 2025; Ni et al., 2023b; Safdar et al., 2025a; Sayagh and Ghafari, 2025; Simoni et al., 2025; Steenhoek et al., 2025a; Sun et al., 2025b; Ullah et al., 2024; Wang et al., 2025a, b, e; Wei et al., 2025; Wen et al., 2025; Weyssow et al., 2025; Widyasari et al., 2024; Yang et al., 2025b, c; Zhu et al., 2025a; Zibaeirad and Vieira, 2025; Chen et al., 2025a; Nguyen et al., 2025a). However, the reliability of LLM-generated reasoning remains under scrutiny. For instance, Steenhoek et al. (Steenhoek et al., 2025a) perform an error analysis on LLM responses, demonstrating that the models struggle to reason about the code semantics relevant to identifying vulnerabilities, especially subtle semantic differences caused by small textual changes. Wen et al. (Wen et al., 2025) propose forward and backward reasoning, where forward reasoning aims to deduce the causes of vulnerabilities, while backward reasoning seeks to understand the code changes in the fixes. Their results show that LLMs can reason about vulnerabilities but often lack the domain-specific knowledge required to effectively invoke such reasoning.
Other approaches request a Structured Vulnerability Report (F2.3) (Dubniczky et al., 2025; Gajjar et al., 2025b; Gnieciak and Szandala, 2025; Keltek et al., 2024; Khare et al., 2025; Mahmud et al., 2025; Noever, 2023; Pelofske et al., 2024; Peng et al., 2025; Shashwat et al., 2024; Sun et al., 2025b; Wei et al., 2025; Weyssow et al., 2025; Widyasari et al., 2025; Xia et al., 2025; Yang et al., 2025c; Yu et al., 2025; Zhou et al., 2025b; Zhu et al., 2025a; Zibaeirad and Vieira, 2025; Jiao et al., 2025; Tsai et al., 2025b; Nguyen et al., 2025a): Studies often aggregate multiple predictive sub-tasks, such as vulnerability localization, severity estimation, type justification, and repair fixes, into a single structured output (e.g., JSON, structured text, or standardized static analysis formats (Gnieciak and Szandala, 2025)). For example, Zibaeirad and Vieira (Zibaeirad and Vieira, 2025) prompt models to generate a report containing the CWE-ID, fix suggestion, vulnerability severity, and exploitability. Similarly, Yang et al. (Yang et al., 2025c) structure the output to include the label, confidence score, vulnerability type, and affected components. Zhou et al. (Zhou et al., 2025b) produce reports detailing function names, namespaces, source/sink type, and dangerous parameter indices.